OpenStudy (anonymous):

Tickets to the movies cost $12 for adults and $8 for children. A group of 14 people went to the movies, and the tickets cost $136. The system of equations models this situation, where x is the number of adults and y is the number of children. How many adults and how many children were in the group? A. 9 adults and 5 children B. 6 adults and 8 children C. 4 adults and 10 children D. 8 adults and 6 children @ganeshie8 @hartnn @mathstudent55 @Abhisar @dan815 @iambatman @Callisto @.Sam. @sangya21

OpenStudy (anonymous):

x+y=14 12x+8y=136

OpenStudy (anonymous):

Correct. So whats x and y?

OpenStudy (anonymous):

idk that was the equation that came with the problem

OpenStudy (anonymous):

x = 14 - y now put this eq in second eq. and find the value of y. then use that value in above eq to find x

OpenStudy (triciaal):

@igivemedalsandfollow! do you understand how they got those equation from the information presented ?

OpenStudy (anonymous):

kinda

OpenStudy (triciaal):

to use the substitution method to solve follow the directions given by @sangya21

OpenStudy (anonymous):

is it c??

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

No. Sorry its not c You can put the values from option in (2) eq.

OpenStudy (anonymous):

its b??

OpenStudy (anonymous):

Correct :)